OpenStudy (anonymous):

What is the sum of the geometric sequence 1, 3, 9, … if there are 12 terms?

OpenStudy (anonymous):

r = 3 a = 1 r>1

OpenStudy (anonymous):

sum= a{r^n - 1}/r-1

OpenStudy (amistre64):

if we construct a general setup of a geo seq: a1 = a1 a2 = a1*r a3 = a1 * r * r a4 = a1 * r * r * r; of a1*r^3 a5 = a1* r^6 a6 = a1*r^5 an = a1*r^(n-1) if we add up all these terms we get: \[S = a1*r^0 + a1*r^1 + a1*r^2 + ... + a1*r^{n-3} + a1*r^{n-2}+ a1*r^{n-1}\] factor out the a1a \[a1*S = r^0 + r^1 + r^2 + ... + r^{n-3} + r^{n-2}+ r^{n-1}\] the trick left is to multiply S by r to get \[S*r = r^1 + r^2 + ... + r^{n-3} + r^{n-2}+ r^{n-1}+r^{n}\] we then subtract one from the other \[\hspace{5em} \ \ S = r^0 + r^1 + r^2 + ... + r^{n-3} + r^{n-2}+ r^{n-1}\]\[\hspace{4em}-(S*r) = -r^1 - r^2 - ... - r^{n-3} - r^{n-2}- r^{n-1}-r^{n}\]-----------------------------------------------------------\[S-(S*r) = r^0 \hspace{16em} - r^{n}\] that was fun .... now lets see; \[S -(S)*r=S(1-r)\] divide both sides by (1-r)\[S=\frac{1-r^n}{1-r}\] \[S=\frac{1-r^n}{1-r}\] distribute the a1 back thru the summation to get\[S=a1\frac{1-r^n}{1-r}\]

OpenStudy (anonymous):

@amistre64 that is only valid when r<1

OpenStudy (anonymous):

here r>1

OpenStudy (amistre64):

since we are not taking a limit to infinity; the summation still holds it is when we want to know if the sum of an infinite geometric sequence converges that we have to consider the magnitude of r

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%281-%283%29%5E12%29%2F%281-3%29 if i factor out the negative from the bottom and push it thru the top we get your version http://www.wolframalpha.com/input/?i=%28%283%29%5E12-1%29%2F%283-1%29 which is identical ...

OpenStudy (amistre64):

@Yahoo!